If tax is 4 units per quantity, new quantity added is 3 units ,what will be the tax revenue,consumer surplus,producer surplus,total surplus


1
Expert's answer
2021-05-09T14:49:58-0400

The tax is a subsidy in form of tax reduction because it increases quantity supplied. In this case price paid by consumers decrease as price received by producers increase as shown in the graph below:





Po=equilibrium price

Qo=equilibrium quantity

P1=price Consumers pay after tax

P2=price producers receive after tax

Q2=quantity supplied after tax

Therefore

(a) tax revenue=tax per unit × quantity sold

"=4\u00d7(Q_0+3)"

"=4Q_0+12"

(b) consumer surplus=area C and D

"\frac{1}{2}×sum \space of \space parallel \space sides× height"

"\frac{1}{2}×[Q_0+(Q_0+3)]×2"

"=2Q_0+3"


(c) producer surplus=area A and B

"\frac{1}{2}×sum \space of \space parallel \space sides× height"

"\frac{1}{2}×[Q_0+(Q_0+3)]×2"

"=2Q_0+3"


(d) Total surplus=Consumer surplus + producer surplus+ tax

"=2Q_0+3 +2Q_0+3+4"

"=4Q_0+10"


For example

If po= 20 and Qo =10

Tax revenue"=4\u00d713=52"

Consumer surplus"=\frac{1}{2}×10×13=65"

Producer surplus"=\frac{1}{2}×10×13=65"

Total surplus"=65+65+4=134"